Express the confidence interval ( 18.7 % , 30.5 % ) ( 18.7 % , 30.5 % ) in the form of ˆ p ± E p ^ ± E . % ± ± %
Answer #1

Since the confidence interval is given as:

( 18.7 % , 30.5 % ) hence

Lower limit =18.7 %

Upper limit=30.5 %

Percentage is calculated as:

percentage=  ( 18.7 %+ 30.5 % )/2

=24.6

so, Margin of Error, E= 24.6-18.7=5.9 or 30.5-24.6=5.9

So, ˆp ± E =24.6% ±5.9%